OpenStudy (anonymous):

A function is defined by Ft=( 0, when - pi

OpenStudy (unklerhaukus):

\[F(t+2\pi)=F(t)=\begin{cases}0,&-\pi<t<-\tfrac\pi2\\2,&-\tfrac\pi2<t<\tfrac\pi2\\0,&~~~\tfrac\pi2<t<\pi\end{cases}\]

OpenStudy (anonymous):

yes thats correct

OpenStudy (unklerhaukus):